OpenStudy (anonymous):

Find the vertex, focus, directrix, and focal width of the parabola. x = 3y2

OpenStudy (welshfella):

the general form of this type of parabola is y^2 = 4ax where the focus is at (a,0) and the directrix is x = -a

OpenStudy (welshfella):

your equation can be written as y^2 = (1/3)x

OpenStudy (welshfella):

so comparing this with the general for we have 4a = 1/3 do you understand this?

OpenStudy (anonymous):

no ;(

OpenStudy (welshfella):

what we are doing is comparing the 2 equations